OpenStudy (jack1):

hey guys, have forgotten how to do partial derivatives, can someone walk me through it plz? have 10 q's, if someone explains one i'm sure i can remember and do the rest

hartnn (hartnn):

sure, ask :)

OpenStudy (jack1):

f(x,y)=y^3x^2 calculate df/dx (x,y)? cheers @hartnn ;)

hartnn (hartnn):

so you are calculating the derivative w.r.t 'x' treat 'y' as constant

hartnn (hartnn):

its like differentiating a^3x^2 or just cx^2

OpenStudy (jack1):

ahh, gotcha, so f(x,y)=y^3x^2 = (k) x^2 so f'(x,y) wrt x = 2k x? = 2 (y^3) x??

hartnn (hartnn):

yes!

OpenStudy (jack1):

sweet, all over this puppy now! cheers @hartnn !!